### 11 Mar 2026 v2

# Geometric masking in AGN jets and its implications for unification and blazar physics

Abstract
We examine the implications of the recently proposed Geometric Masking scenario for the Active Galactic Nuclei (AGN) Unification Scheme, the blazar sequence, and broader blazar phenomenology, and evaluate its consistency with recent observations. Phase-resolved analysis of PKS 2155-304 indicates that extreme GeV spectral hardening events are locked to the trough of its $\sim 1.7$ yr quasi-periodic oscillation, suggesting that high-flux states are dominated by a soft, geometrically amplified envelope that outshines an underlying hard core. Independent support comes from the detection of TeV emission in the FSRQs S5 1027+74 and 3C 273 when integrating over low-flux states, revealing hard spectra that are difficult to reconcile with a purely cooling-dominated interpretation. We suggest that Doppler boosting preferentially enhances the soft component when the jet is closely aligned, creating a visibility bias that extends the AGN Unification Scheme and introduces a geometric modulation layer within the blazar sequence. In this framework, low-flux states correspond to windows of geometric transparency rather than intrinsic quiescence, and misaligned systems act as largely unmasked laboratories of particle acceleration. The scenario indicates that the duty cycle of extreme acceleration in AGN jets may be higher than inferred from flux-selected observations.
Active galactic nuclei (16) — Gamma-ray sources (633) — Radio jets (1347)
I 1. Introduction
Blazar gamma-ray emission is largely dominated by stochastic variability, commonly modeled as red-noise processes in their power spectral densities (e.g., Abdo et al., 2010; Vaughan et al., 2016). However, a minority of sources show year-scale quasi-periodic oscillations (QPOs), whose origin remains debated between geometric mechanisms such as jet precession and intrinsic plasma-driven instabilities (e.g., Ackermann et al., 2015; Peñil et al., 2020; Rico et al., 2025). Because timing information alone is insufficient to discriminate between these scenarios, phase-resolved spectral analyses provide an additional diagnostic (Madero & Domínguez, 2026).
The Active Galactic Nuclei (AGN) Unification Scheme explains the diversity of AGN classes primarily through orientation and relativistic beaming (e.g., Urry & Padovani, 1995), while the blazar sequence relates luminosity and spectral properties to intrinsic jet power and radiative cooling (Fossati et al., 1998; Ghisellini et al., 1998). Although often discussed separately, both geometry and intrinsic physics contribute to the observed phenomenology.
In a recent study of the high-synchrotron peak (HSP) blazar PKS 2155-304, we reported a GeV spectral hardening event, manifested as a high-energy upturn, phase-locked to the trough of its $\sim 1.7$ yr QPO (Dinesh et al., 2025; Domínguez et al., 2026). Similarly, the periodic blazar PG 1553+113 ($P\approx 2.2$ yr) shows a global softer-when-brighter trend, differing from the standard shock-in-jet expectation of spectral hardening during flux peaks (e.g., Ackermann et al., 2015; Madero & Domínguez, 2026).
These results suggest a Geometric Masking scenario in which high-flux states are dominated by a soft, geometrically boosted envelope (the mask), while plasma-driven hard-spectrum signatures (the core) become more observable when geometric boosting is reduced (Domínguez et al., 2026). This visibility effect may operate within the AGN Unification Scheme and introduce a state-dependent modulation of spectral contrast, and it may also imply that part of the diversity associated with the blazar sequence arises from Doppler-dependent visibility effects. Independent support comes from the TeV detection of the FSRQs S5 1027+74 (Paliya et al., 2025) and 3C 273 (VERITAS Collaboration, 2026) during integrated low-flux states, revealing hard spectral components that are consistent with relativistic masking affecting the detectability of extreme particle acceleration. In this Research Note, we examine the implications of this hypothesis for AGN unification, the blazar sequence, and broader blazar phenomenology, including misaligned sources and acceleration duty cycles.
II 2. The Physical Mechanism
The Geometric Masking scenario arises from the interplay between relativistic beaming and the spectral properties of distinct jet components. We consider a two-component structure: (1) a soft-spectrum Mask (sheath/envelope), such as emission from cooled electrons, strongly modulated by geometry; and (2) a hard-spectrum Core (spine), stochastic and associated with fresh particle acceleration at shocks, either less strongly beamed or intrinsically fainter than the boosted mask.
The visibility of the Core is regulated by the Doppler factor, $\delta=[\Gamma(1-\beta\cos\theta)]^{-1}$. The observed flux density scales as $F_{\nu}\propto\delta^{3+\alpha}$, with spectral index $\alpha$ ($F_{\nu}\propto\nu^{-\alpha}$). The soft Mask typically has a steeper index ($\alpha_{\rm mask}\sim 1.5$) than the hard Core ($\alpha_{\rm core}\sim 0.5$).
As the viewing angle $\theta$ decreases and $\delta$ increases, the soft Mask is amplified more rapidly ($\propto\delta^{4.5}$) than the hard Core ($\propto\delta^{3.5}$). In high-flux aligned states, this differential boosting can allow the soft emission to dominate, effectively masking the central engine. During geometric minima (QPO troughs or misaligned views), $\delta$ is reduced, dimming the Mask and increasing the contrast of the Core, thereby making spectral hardening events more observable (Domínguez et al., 2026).
This mechanism provides a possible explanation for the softer-when-brighter behavior observed in PKS 2155-304 (Domínguez et al., 2026) and PG 1553+113 (Madero & Domínguez, 2026), where flux peaks are dominated by boosted soft emission rather than fresh particle injection.
III 3. A Visibility-Regulated Framework for AGN and Blazars
III.1 3.1. Implications for the AGN Unification Scheme
The Geometric Masking scenario posits that the visibility of the central engine is regulated by the Doppler factor ($\delta$). If high-$\delta$ states amplify the soft component more efficiently than the hard stochastic core, a visibility bias may arise with consequences for AGN unification, the blazar sequence, and broader blazar phenomenology.
Within the AGN Unification Scheme, orientation determines whether a source appears as a blazar or a misaligned radio galaxy. Our results extend this framework by indicating that orientation and Doppler boosting regulate not only class identity, but also which spectral component dominates at a given time. Unification therefore functions as both a classification scheme and a visibility scheme.
Standard unification models associate FSRQs with soft gamma-ray spectra due to cooling in the broad line region. In this framework, hard intrinsic acceleration signatures may be concealed beneath boosted soft emission. Support comes from TeV detections of FSRQs S5 1027+74 (Paliya et al., 2025) and 3C 273 (VERITAS Collaboration, 2026), obtained by integrating over low-flux states rather than flares, revealing hard spectral components. In this scenario, low-flux states correspond to transparent windows where the geometric mask dims and exposes the hard spine, suggesting efficient particle acceleration may be common in FSRQ jets but often hidden by relativistic amplification.
III.2 3.2. Implications for the Blazar Sequence
For the blazar sequence, part of the observed luminosity–spectral trends may reflect contrast and selection effects in addition to intrinsic jet power and cooling physics. If increasing $\delta$ preferentially boosts the soft component, the most aligned sources may appear more luminous and spectrally softer, whereas low-$\delta$ views or geometric minima reveal harder components at lower apparent luminosity. Intrinsic drivers such as accretion regime and radiative cooling remain fundamental, but geometry introduces a modulation layer that can shift individual sources across the sequence plane between masked (high-$\delta$) and unmasked (low-$\delta$) states.
Our results favor a two-component jet structure comprising a soft, geometrically modulated envelope (sheath) and a stochastic, hard-spectrum core (spine) (Domínguez et al., 2026). The distinction between FSRQs and HSPs during flares may therefore be partly determined by which component dominates geometrically at the observer’s viewing angle. In closely aligned systems, the mask can obscure the spine, suggesting that the intrinsic duty cycle of extreme particle acceleration across the AGN population may be higher than current surveys indicate.
Furthermore, the Geometric Masking scenario provides a possible explanation for the observation that FSRQs are significantly more variable in gamma rays than BL Lacs. In FSRQs, dense external photon fields from the broad-line region or torus act as an amplifier: intrinsic acceleration events scatter these photons via External Compton, producing highly luminous flares that can overpower the steady geometric mask. In HSPs lacking these external fields, flares rely on the less efficient Synchrotron Self-Compton process. These intrinsically weaker flares may remain observationally dominated by the geometrically boosted baseline envelope, becoming more visible when the Doppler mask is minimized. Consequently, the intrinsic duty cycle of extreme particle acceleration may be comparable across both classes, but it can be more difficult to observe in HSPs.
III.3 3.3. Broader Blazar Phenomenology
If aligned sources are mask-dominated due to high Doppler boosting, misaligned sources (radio galaxies) may be effectively permanently unmasked. Systems such as Cen A or M87, with lower Doppler factors, reduce amplification of the soft sheath relative to the hard core (assuming $\alpha_{\rm soft}>\alpha_{\rm hard}$). Although flux-limited, these sources are contrast-rich, and intrinsic hard spectral features may appear with a higher duty cycle than in blazars, since geometric dips are not required to enhance contrast. Observational strategies may therefore benefit from prioritizing deep spectral integration of misaligned sources to probe acceleration physics with reduced geometric masking.
This interpretation differs from the standard brighter-is-better strategy and addresses a tension with conventional jet models, where a faster spine ($\Gamma_{\rm spine}>\Gamma_{\rm sheath}$) would predict harder-when-brighter behavior. Doppler boosting scales as $\delta^{3+\alpha}$ (e.g., Urry & Padovani, 1995). Because the soft mask typically has a steeper spectrum ($\alpha_{\rm soft}\sim 1.5$) than the hard core ($\alpha_{\rm hard}\sim 0.5$), increasing $\delta$ amplifies the soft component more rapidly ($\propto\delta^{4.5}$) than the hard one ($\propto\delta^{3.5}$), allowing soft emission to dominate in high-Doppler states. This is consistent with the observed softer-when-brighter trends (Domínguez et al., 2026; Madero & Domínguez, 2026) and does not exclude rare extreme outbursts, in which a compact hard-emitting region may outshine the boosted soft sheath and temporarily overcome the masking effect.
Finally, low-flux states are interpreted as intrinsic quiescence (“engine off”), but in this framework they represent windows of geometric transparency. The detection of hard spectral components during flux minima, both in the periodic trough of PKS 2155-304 (Domínguez et al., 2026) and in the integrated low-flux states of S5 1027+74 and 3C 273 (Paliya et al., 2025; VERITAS Collaboration, 2026), challenges this interpretation and suggests that the intrinsic duty cycle of particle acceleration may be high, while the observational duty cycle is regulated by the masking effect.
IV 4. Conclusion
We suggest that the AGN Unification Scheme can be viewed as a visibility-regulated framework in which orientation and Doppler boosting determine not only class appearance but also the contrast between jet components. The Geometric Masking scenario, supported by the phase-locked spectral hardening in PKS 2155-304 (Domínguez et al., 2026) and the hard-spectrum detections of S5 1027+74 and 3C 273 during low-flux states (Paliya et al., 2025; VERITAS Collaboration, 2026), indicates that high-flux states are often dominated by differentially boosted soft emission. Broader observational tests are needed to assess the robustness and generality of this interpretation.
In this framework, what is traditionally termed quiescence in aligned jets corresponds to a window of geometric transparency, while misaligned systems act as naturally unmasked laboratories where the hard core remains visible. The blazar sequence therefore reflects both intrinsic jet physics and geometry-driven visibility effects, and the intrinsic duty cycle of extreme particle acceleration may be higher than inferred from flux-selected surveys.
To further probe the physics of the central engine, future high-sensitivity surveys could combine orientation-aware strategies with deep observations of geometric minima, where the mask is faintest and the core most exposed.